planting grass on the rectangular plot of land and also building a fence around the edge of the plot . The plot is 45 yards long and 30 yards wide . how much area do we need to cover with grass seed ? how many feet of fencing do we need ?
Answer by nerdybill(7384)

planting grass on the rectangular plot of land and also building a fence around the edge of the plot . The plot is 45 yards long and 30 yards wide . how much area do we need to cover with grass seed ? how many feet of fencing do we need ?
.
Convert yards to feet so:
45 yards long and 30 yards wide
is the same as:
135 feet long and 90 feet wide
.
how much area do we need to cover with grass seed?
135*90 = 12150 square feet
.
how many feet of fencing do we need ?
2(135+90) = 2(225) = 450 feet